Ten years ago, an account was opened with an investment of $5000. Its balance is now $8000. If interest was compounded annually, what interest rate was earned on the account? (Enter your answers as a decimal rounded to 4 decimal places, not a percentage. For example, enter 0.0843 instead of 8.43%)
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

8000=5000*(1+r/100)^10

(8000/5000)^(1/10)=(1+r/100)

(1+r/100)=1.0481

r=1.0481-1

=0.0481(Approx).